# On the variety of general position problems under vertex and edge removal

### v1

@@ 1 Introduction
The concept of a general position set was introduced into graph theory independently in two papers, namely in [4, 12], but in the case of hypercubes implicitly also much earlier in [9]. It was the article [12] that has stimulated a great deal of interest in this concept. The survey [3] provides a systematic review of the results, in addition we point to the following selected recent publications [2, 5, 8, 10, 11, 14, 15, 16, 19].
For a general position set, one requires that the vertices of the set are in general position. This requirement can be generalized by requiring that the vertices of some set lie in general position with respect to the selected set. Using this idea, and considering the natural selections for sets to be considered, in [17] a variety of general position sets was defined as follows. First, for a given graph $G=(V(G),E(G))$ and a set of vertices $Z\subseteq V(G)$, two vertices $x,y\in V(G)$ are said to be $Z$-positionable if any shortest $x,y$-path intersects $Z$ at most in $x$ and $y$, that is, no inner vertex of the path lies in $Z$. Second, if $\overline{Z}=V(G)\setminus Z$, then the set $Z$ is
@@
Typically, we are interested in the cardinalities of largest such sets, which are respectively denoted by ${\rm gp}(G)$, ${\rm gp}_{\rm t}(G)$, ${\rm gp}_{\rm o}(G)$, and ${\rm gp}_{\rm d}(G)$, and called the general position number, the total general position number, the outer general position number, and the dual general position number of $G$. If $\tau\in\{{\rm gp},{\rm gp}_{\rm t},{\rm gp}_{\rm o},{\rm gp}_{\rm d}\}$, then a set $Z$ is a $\tau$-set if it has the corresponding property and $|Z|=\tau(G)$. Note that, by definition, ${\rm gp}_{\rm t}(G)\leq{\rm gp}_{\rm o}(G)\leq{\rm gp}(G)$ and ${\rm gp}_{\rm t}(G)\leq{\rm gp}_{\rm d}(G)\leq{\rm gp}(G)$. After being introduced in [17], this variety has been also investigated on strong and lexicographic products of graphs [6].
@@
In the next section we give definitions and recall results needed in the rest of the paper. In Section 3 we consider the vertex removal. For the total general position sets we prove that if $x$ is not a cut vertex, then ${\rm gp}_{\rm t}(G)-1\leq{\rm gp}_{\rm t}(G-x)\leq{\rm gp}_{\rm t}(G)+\deg_{G}(x)$. Section 3.2 addresses the outer general position number and informs that anything can happen with outer general position sets when a vertex is removed. On the other hand, if $x$ lies in some ${\rm gp}_{\rm o}$-set, then ${\rm gp}_{\rm o}(G)-1\leq{\rm gp}_{\rm o}(G-x)$. Moreover, if $x$ is a simplicial vertex, then ${\rm gp}_{\rm o}(G-x)\leq{\rm gp}_{\rm o}(G)+\deg_{G}(x)-1$. Section 3.3 then deals with the dual general position number. Similarly, as for the outer general position number, ${\rm gp}_{\rm d}(G-x)$ can be arbitrarily larger/smaller than ${\rm gp}_{\rm d}(G)$. On the other hand, if $x$ is not a cut vertex and lies in some ${\rm gp}_{\rm d}$-set of $G$, then ${\rm gp}_{\rm d}(G)-1\leq{\rm gp}_{\rm d}(G-x)$. In the final section we consider the edge removal operation. For the total and the outer general position number we prove sharp lower and upper bounds, while for the dual general position number we show that the difference ${\rm gp}_{\rm d}(G)-{\rm gp}_{\rm d}(G-e)$ can be arbitrarily large. All bounds proved in this paper are demonstrated to be sharp.
@@ 2 Preliminaries
Let $u\in V(G)$. Then $N_{G}(u)$ denotes the set of neighbors of $u$ in $G$ and $N_{G}[u]=N_{G}(u)\cup\{u\}$. The degree $\deg_{G}(u)$ of $u$ is $\deg_{G}(u)=|N_{G}(u)|$. If $X\subseteq V(G)$, then the subgraph of $G$ induced by $X$ is denoted by $G[X]$. The graph GuG-u is obtained by deleting $u$ and all incident edges from $G$, that is, GuG-u is the subgraph $G[V(G)\setminus\{u\}]$. For an edge $e\in E(G)$, the graph GeG-e is obtained by deleting the edge $e$ from $G$. A vertex $u$ of $G$ is simplicial if $N_{G}(u)$ induces a complete subgraph. The set of simplicial vertices of $G$ will be denoted by $S(G)$ and the cardinality of $S(G)$ by $s(G)$. Moreover, $\omega(G)$ and $\alpha(G)$ stand for the clique number and the independence number of $G$.
The distance $d_{G}(u,v)$ between vertices $u$ and $v$ of $G$ is the number of edges on a shortest $u,v$-path. Let $H$ be a subgraph of $G$. Then $H$ is convex if for any vertices $u,v\in V(H)$, any shortest $u,v$-path in $G$ lies completely in $H$. By abuse of language, we will also say that a set $X\subseteq V(G)$ is convex if $X$ induces a convex subgraph.
The vertex $u$ of $G$ is maximally distant from a vertex $v\in V(G)$ if every $w\in N_{G}(u)$ satisfies $d_{G}(v,w)\leq d_{G}(u,v)$. If $u$ is maximally distant from $v$, and $v$ is also maximally distant from $u$, then $u$ and $v$ are mutually maximally distant, MMD for short. Note that true twins, that is, vertices $u$ and $v$ with $N_{G}[u]=N_{G}[v]$, are MMD. The strong resolving graph $G{}_{\rm SR}$ of $G$ has $V(G)$ as the vertex set, two vertices being adjacent in $G{}_{\rm SR}$ if they are MMD in $G$. This concept was introduced in [13] for the purpose of better understanding the strong metric dimension of graphs.
The first known result which we need later on describes general position sets in an arbitrary graph. To state it, some more definitions are required. If ${\cal P}=\{X_{1},\ldots,X_{t}\}$ a partition of $X\subseteq V(G)$, then $\cal P$} is distance-constant if for any $\in[t], iji\neq j$, there exists a constant $p_{ij}$, such that $d_{G}(x,y)=p_{ij}$ for every $x\in X_{i}$, $y\in X_{j}$. If so, we set $d_{G}(X_{i},X_{j})=p_{ij}$. A distance-constant partition $\cal P$} is in-transitive if $p_{ik}\neq p_{ij}+p_{jk}$ holds for $i,j,k\in[t]$.
@@
The following theorem is the most important for us since it contains characterizations of the three variants of general position sets. They were respectively proved in [17, Theorems 2.1, 2.3, 3.1].
@@ 3 Vertex removal
Proof. Since $s(G_{n})=2n$, Theorem 2.2(i) yields ${\rm gp}_{\rm t}(G_{n})=2n$. Observe that two vertices of $G_{n}$ are MMD if and only if they are leaves. Hence ${\rm gp}_{\rm o}(G_{n})=2n$ by Theorem 2.2(ii). By Lemma 2.3 we also have ${\rm gp}_{\rm d}(G_{n})\geq 2n$. To prove the reverse inequality, we will apply Theorem 2.2(iii). For this sake we first get by a case analysis that ${\rm gp}(G_{n})=2n+1$. By symmetry, the only $\rm gp$}-sets to be considered are $X_{1}\cup X_{2}\cup\{v\}$, $X_{2}\cup Y\cup\{u\}$, $X_{2}\cup Y\cup\{v\}$, and $X_{2}\cup Y\cup\{w\}$. If $Z$ is any of these sets, then $G_{n}-Z$ is not connected, hence clearly not convex. Theorem 2.2(iii) thus implies that none of these sets is a dual general position set. It follows that ${\rm gp}_{\rm d}(G_{n})\leq 2n$ and we can conclude that ${\rm gp}_{\rm d}(G_{n})=2n$.
@@ 3.1 Total general position sets
In this subsection we show that the total general position set of GxG-x can be bounded below and above when $x$ is not a cut vertex of graph $G$.
@@
To prove the lower bound it suffices to show that $X\setminus\{x\}\subseteq S(G-x)$. Suppose on the contrary that $y\in X\setminus\{x\}$ is not simplicial in GxG-x. Then there exist $z,w\in N_{G-x}(y)$ such that $z$ and $w$ are not adjacent in GxG-x. The vertices $z$ and $w$ are also not adjacent in $G$, but then $y$ is not simplicial in $G$, a contradiction. Hence $X\setminus\{x\}\subseteq S(G-x)$ and by applying Theorem 2.2(i) we get ${\rm gp}_{\rm t}(G)-1\leq{\rm gp}_{\rm t}(G-x)$. This proves the lower bound.
@@ 3.2 Outer general position sets
To show the bound, it suffices to prove the assertion that $X\setminus\{x\}$ is an outer general position set of GxG-x. Given Theorem 2.2(ii) this is equivalent to proving that every two vertices in $X\setminus\{x\}$ are MMD in GxG-x. Suppose on the contrary that there exist $u,v\in X\setminus\{x\}$ such that $u$ and $v$ are not MMD in GxG-x. By symmetry we may assume that there exists $u^{\prime}\in N_{G-x}(u)$ such that $d_{G-x}(u^{\prime},v)=d_{G-x}(u,v)+1$. Since $u$ and $v$ are MMD in $G$, we have $d_{G}(u^{\prime},v)\leq d_{G}(u,v)$. This implies that $x$ lies on some shortest $u^{\prime},v$-path in $G$. But then $v$ and $x$ are not MMD in $G$, a contradiction. This contradiction implies that every pair of vertices in $X\setminus\{x\}$ are MMD in GxG-x. We can conclude that ${\rm gp}_{\rm o}(G-x)\geq|X\setminus\{x\}|={\rm gp}_{\rm o}(G)-1$.
@@ 3.3 Dual general position sets
Consider the fan graphs $F_{n}$, $n\geq 4$, with the vertex $x$ of degree $n$. By Theorem 2.2(iii), it is straightforward to check that every largest general position set of $F_{n}$ is also a dual general position set. Hence ${\rm gp}(F_{n})={\rm gp}_{\rm d}(F_{n})$. Moreover, it was proved in [18, Corollary 2.9] that if $n\geq 4$, then ${\rm gp}(F_{n})=\left\lceil\frac{2(n+1)}{3}\right\rceil$. Hence, since $F_{n}-x\cong P_{n}$, we have
${\rm gp}_{\rm d}(F_{n})=\left\lceil\frac{2(n+1)}{3}\right\rceil\quad{\rm and}\quad{\rm gp}_{\rm d}(F_{n}-x)=2\,.$
@@
To prove that ${\rm gp}_{\rm d}(M_{k})\leq k+2$, suppose on the contrary that there exists a dual general position set $X$ of $M_{k}$ with $|X|>k+2$. We first claim that $x\notin X$. Suppose that $x\in X$. Then at least one of $w_{1}$ and $w_{4}$ must belong to $X$, say $w_{1}\in X$. Then $w_{4}\notin X$. Since $w_{1}\in X$, exactly one of $w_{2}$ and $w_{k+4}$ lies in $X$. If $w_{k+4}\in X$, then $w_{2}$ and $w_{4}$ are not $X$-positionable. And if $w_{2}\in X$, then $w_{k+4}$ and $w_{4}$ are not $X$-positionable. We analogously get a contradiction in the case when $w_{4}\in X$. We can conclude that $x\notin X$. We next claim that $w_{i}\notin X$ for each $i\in[k]$. Suppose that $w_{i}\in X$ for some $i\in[k]$. Since $x\notin X$, we get that $v_{i}\in X$ because $x,w_{i},v_{i}$ is the unique shortest $x,v_{i}$-path. Further, exactly one of the neighbors of $w_{i}$ on the $(k+4)$-cycle of $W_{k+4}$ lies in $X$. But this is not possible, as $X$ is a general position set.
@@
Proof. Let $X$ be a ${\rm gp}_{\rm d}$-set of $G$ such that $x\in X$. Then $|X|={\rm gp}_{\rm d}(G)$ and by Theorem 2.2(ii), GXG-X is convex. By the proof of [7, Proposition 3.3], the set $X\setminus\{x\}$ is a general position set of GxG-x. It suffices to show that $(G-x)-(X\setminus\{x\})$ is convex in GxG-x. Suppose on the contrary that this is not the case. There are $u,v\notin X$ such that a shortest $u,v$-path $P$ containing a vertex $w\in X\setminus\{x\}$. Since GXG-X is convex, $d_{G}(u,v)<d_{G}(u,w)+d_{G}(w,v)$. Based on this, we can conclude the following:
@@
This in turn implies that $x$ lies on the shortest $u,v$-path in $G$ which contradicts with GXG-X being convex. We can conclude that $(G-x)-(X\setminus\{x\})$ is convex in GxG-x. Using Theorem 2.2(iii) again, the set $X\setminus\{x\}$ is a dual general position set of GxG-x and hence ${\rm gp}_{\rm d}(G-x)\geq|X\setminus\{x\}|={\rm gp}_{\rm d}(G)-1$.
@@
To prove ${\rm gp}_{\rm d}(Y_{k})\leq 5$, consider an arbitrary ${\rm gp}_{\rm d}$-set $X$ of $Y_{k}$. Suppose on the contrary that $v_{i}\in X$ for some $i\in[k]$. Then exactly one of the vertices $x$ and $y$ must be in $X$. Otherwise, $v_{i}$ would lie on a shortest $x,y$-path, which is a contradiction. Assume, without loss of generality, that $x\in X$ and $y\notin X$. We claim that there is no vertex from $\{u,v,w\}$ contained in $X$. If $z\in\{u,v,w\}\cap X$, then $v_{i}$ lies on a shortest $x,z$-path, contradicts with the fact that $X$ is a ${\rm gp}_{\rm d}$-set of $Y_{k}$. If a vertex of $\{u,v,w\}$ does not contain in $X$, say $u$, then $v_{i}$ lies on a shortest $y,u$-path, which implies that $y$ and $u$ are not $X$-positionable. This contradiction yields $v_{i}\notin X$ for each $i\in[k]$. Therefore, $|X|\leq 5$, and we conclude that ${\rm gp}_{\rm d}(Y_{k})=5$.

### v2

@@ 1 Introduction
The concept of a general position set was introduced into graph theory independently in two papers, namely in [4, 13], but in the case of hypercubes, implicitly also much earlier in [10]. It was the article [13] that has stimulated a great deal of interest in this concept. The survey [3] provides a systematic review of the results, in addition we point to the following selected recent publications [2, 5, 8, 9, 11, 12, 15, 16, 17, 20].
For a general position set, that is, a set in which no three vertices lie on a common shortest path, one requires that the vertices of the set are in general position. This requirement can be generalized by requiring that the vertices of some set lie in general position with respect to the selected set. Using this idea, and considering the natural selections for sets to be considered, in [18] a variety of general position sets was defined as follows. First, for a given graph $G=(V(G),E(G))$ and a set of vertices $Z\subseteq V(G)$, two vertices $x,y\in V(G)$ are said to be $Z$-positionable if any shortest $x,y$-path intersects $Z$ at most in $x$ and $y$, that is, no inner vertex of the path lies in $Z$. Second, if $\overline{Z}=V(G)\setminus Z$, then the set $Z$ is
@@
Typically, we are interested in the cardinalities of largest such sets, which are respectively denoted by ${\rm gp}(G)$, ${\rm gp}_{\rm t}(G)$, ${\rm gp}_{\rm o}(G)$, and ${\rm gp}_{\rm d}(G)$, and called the general position number, the total general position number, the outer general position number, and the dual general position number of $G$, respectively. If $\tau\in\{{\rm gp},{\rm gp}_{\rm t},{\rm gp}_{\rm o},{\rm gp}_{\rm d}\}$, then a set $Z$ is a $\tau$-set if it has the corresponding property and $|Z|=\tau(G)$. Note that, by definition, ${\rm gp}_{\rm t}(G)\leq{\rm gp}_{\rm o}(G)\leq{\rm gp}(G)$ and ${\rm gp}_{\rm t}(G)\leq{\rm gp}_{\rm d}(G)\leq{\rm gp}(G)$. After being introduced in [18], this variety has been also investigated on strong and lexicographic products of graphs [6].
@@
In the next section we give definitions and recall results needed in the rest of the paper. In Section 3 we consider the vertex removal. For the total general position sets we prove that if $x$ is not a cut vertex, then ${\rm gp}_{\rm t}(G)-1\leq{\rm gp}_{\rm t}(G-x)\leq{\rm gp}_{\rm t}(G)+\deg_{G}(x)$. Section 3.2 addresses the outer general position number of a graph when a vertex is removed and shows that it cannot be bounded in terms of the outer general position number of the original graph. On the other hand, if $x$ lies in some ${\rm gp}_{\rm o}$-set, then ${\rm gp}_{\rm o}(G)-1\leq{\rm gp}_{\rm o}(G-x)$. Moreover, if $x$ is a simplicial vertex, then ${\rm gp}_{\rm o}(G-x)\leq{\rm gp}_{\rm o}(G)+\deg_{G}(x)-1$. Section 3.3 then deals with the dual general position number. Similarly, as for the outer general position number, ${\rm gp}_{\rm d}(G-x)$ can be arbitrarily larger/smaller than ${\rm gp}_{\rm d}(G)$. On the other hand, if $x$ is not a cut vertex and lies in some ${\rm gp}_{\rm d}$-set of $G$, then ${\rm gp}_{\rm d}(G)-1\leq{\rm gp}_{\rm d}(G-x)$. In the final section we consider the edge removal operation. For the total and the outer general position number we prove sharp lower and upper bounds, while for the dual general position number we show that the difference ${\rm gp}_{\rm d}(G)-{\rm gp}_{\rm d}(G-e)$ can be arbitrarily large. All bounds proved in this paper are demonstrated to be sharp.
@@ 2 Preliminaries
Let $u\in V(G)$. Then $N_{G}(u)$ denotes the set of neighbors of $u$ in $G$ and $N_{G}[u]=N_{G}(u)\cup\{u\}$. The degree $\deg_{G}(u)$ of $u$ is $\deg_{G}(u)=|N_{G}(u)|$. If $X\subseteq V(G)$, then the subgraph of $G$ induced by $X$ is denoted by $G[X]$. The graph GuG-u is obtained by deleting $u$ and all incident edges from $G$, that is, GuG-u is the subgraph $G[V(G)\setminus\{u\}]$. For an edge $e\in E(G)$, the graph GeG-e is obtained by deleting the edge $e$ from $G$. A vertex $u$ of $G$ is simplicial if $N_{G}(u)$ induces a complete subgraph. The set of all simplicial vertices of $G$ will be denoted by $S(G)$ and the cardinality of $S(G)$ by $s(G)$. Moreover, $\omega(G)$ and $\alpha(G)$ stand for the clique number and the independence number of $G$.
The distance $d_{G}(u,v)$ between vertices $u$ and $v$ of $G$ is the number of edges on a shortest $u,v$-path. A set $X\subseteq V(G)$ is convex if for any vertices $u,v\in X$, any shortest $u,v$-path contains only vertices from $X$. By abuse of language, we will also say that a subgraph $H$ of $G$ is convex if $V(H)$ is convex. The vertex $u$ of $G$ is maximally distant from a vertex $v\in V(G)$ if every $w\in N_{G}(u)$ satisfies $d_{G}(v,w)\leq d_{G}(u,v)$. If $u$ is maximally distant from $v$, and $v$ is also maximally distant from $u$, then $u$ and $v$ are mutually maximally distant, MMD for short. Note that true twins, that is, vertices $u$ and $v$ with $N_{G}[u]=N_{G}[v]$, are MMD. The strong resolving graph $G{}_{\rm SR}$ of $G$ has $V(G)$ as the vertex set, two vertices being adjacent in $G{}_{\rm SR}$ if they are MMD in $G$. This concept was introduced in [14] for the purpose of better understanding the strong metric dimension of graphs.
The first known result which we need later on describes general position sets in an arbitrary graph. To state it, some more definitions are required. If ${\cal P}=\{X_{1},\ldots,X_{t}\}$ is a partition of $X\subseteq V(G)$, then $\cal P$} is distance-constant if for any $\in[t], iji\neq j$, there exists a constant $p_{ij}$, such that $d_{G}(x,y)=p_{ij}$ for every $x\in X_{i}$, $y\in X_{j}$. If so, we set $d_{G}(X_{i},X_{j})=p_{ij}$. A distance-constant partition $\cal P$} is in-transitive if $p_{ik}\neq p_{ij}+p_{jk}$ holds for $i,j,k\in[t]$.
@@
The following theorem is the most important for us since it contains characterizations of the three variants of general position sets. They were respectively proved in [18, Theorems 2.1, 2.3, 3.1].
@@ 3 Vertex removal
Proof. Since $s(G_{n})=2n$, Theorem 2.2(i) yields ${\rm gp}_{\rm t}(G_{n})=2n$. Observe that two vertices of $G_{n}$ are MMD if and only if they are leaves, or they both belong to $X_{1}$, or they both belong to $X_{2}$. Hence the strong resolving graph $(G_{n}){}_{\rm SR}$ has three non-trivial components, one isomorphic to $K_{2n}$, and the other two each isomorphic to $K_{n}$. Hence ${\rm gp}_{\rm o}(G_{n})=2n$ by Theorem 2.2(ii). By Lemma 2.3 we also have ${\rm gp}_{\rm d}(G_{n})\geq 2n$. To prove the reverse inequality, we will apply Theorem 2.2(iii). For this sake we first get by a case analysis that ${\rm gp}(G_{n})=2n+1$. By symmetry, the $\rm gp$}-sets to be considered are $X_{1}\cup X_{2}\cup\{v\}$, $X_{2}\cup Y\cup\{u\}$, $X_{2}\cup Y\cup\{v\}$, $X_{2}\cup Y\cup\{w\}$, $X_{1}\cup Y\cup\{u\}$, $X_{1}\cup Y\cup\{v\}$, and $X_{1}\cup Y\cup\{w\}$. If $Z$ is any of these sets, then $G_{n}-Z$ is not connected, hence clearly not convex. Theorem 2.2(iii) thus implies that none of these sets is a dual general position set. It follows that ${\rm gp}_{\rm d}(G_{n})\leq 2n$ and we can conclude that ${\rm gp}_{\rm d}(G_{n})=2n$.
@@ 3.1 Total general position sets
In this subsection we show that the total general position set of GxG-x can be bounded below and above when $x$ is not a cut vertex of a graph $G$.
@@
To prove the lower bound, it suffices to show that $X\setminus\{x\}\subseteq S(G-x)$. Suppose to the contrary that $y\in X\setminus\{x\}$ is not simplicial in GxG-x. Then there exist $z,w\in N_{G-x}(y)$ such that $z$ and $w$ are not adjacent in GxG-x. The vertices $z$ and $w$ are also not adjacent in $G$, but then $y$ is not simplicial in $G$, a contradiction. Hence $X\setminus\{x\}\subseteq S(G-x)$ and by applying Theorem 2.2(i) we get ${\rm gp}_{\rm t}(G)-1\leq|X\setminus\{x\}|\leq{\rm gp}_{\rm t}(G-x)$. This proves the lower bound.
@@
Note that in view of Theorem 2.2(i), we can rephrase Theorem 3.2 by saying that if $x$ is not a cut vertex of a graph $G$, then
$s(G)-1\leq s(G-x)\leq s(G)+\deg_{G}(x)\,$,
and if $x$ is a simplicial vertex, then
$s(G-x)\leq s(G)+\deg_{G}(x)-1\,.$
@@ 3.2 Outer general position sets
To show the bound, it suffices to prove the assertion that $X\setminus\{x\}$ is an outer general position set of GxG-x. Given Theorem 2.2(ii) this is equivalent to proving that every two vertices in $X\setminus\{x\}$ are MMD in GxG-x. Suppose to the contrary that there exist $u,v\in X\setminus\{x\}$ such that $u$ and $v$ are not MMD in GxG-x. By symmetry we may assume that there exists $u^{\prime}\in N_{G-x}(u)$ such that $d_{G-x}(u^{\prime},v)=d_{G-x}(u,v)+1$. Since $u$ and $v$ are MMD in $G$, we have $d_{G}(u^{\prime},v)\leq d_{G}(u,v)$. This implies that $x$ lies on some shortest $u^{\prime},v$-path in $G$. But then $v$ and $x$ are not MMD in $G$, a contradiction. This contradiction implies that every pair of vertices in $X\setminus\{x\}$ are MMD in GxG-x. We can conclude that ${\rm gp}_{\rm o}(G-x)\geq|X\setminus\{x\}|={\rm gp}_{\rm o}(G)-1$.
@@ 3.3 Dual general position sets
Consider the fan graphs $F_{n}$ ($n\geq 4$) defined in Subsection 3.2, where $x$ denotes the vertex of degree $n$. It was proved in [19, Corollary 2.9] that if $n\geq 4$, then ${\rm gp}(F_{n})=\left\lfloor\frac{2(n+1)}{3}\right\rfloor$. Moreover, a gp-set of $F_{n}$ consists of the end-vertices of a largest set of independent edges of the path $P_{n}$ in $F_{n}$. By Theorem 2.2(iii) it is then straightforward to check that every largest general position set of $F_{n}$ is a dual general position set. Hence ${\rm gp}(F_{n})={\rm gp}_{\rm d}(F_{n})$. Moreover, since ${\rm gp}(F_{n})=\left\lfloor\frac{2(n+1)}{3}\right\rfloor$ for $n\geq 4$, and because $F_{n}-x\cong P_{n}$, we have
${\rm gp}_{\rm d}(F_{n})=\left\lfloor\frac{2(n+1)}{3}\right\rfloor\quad{\rm and}\quad{\rm gp}_{\rm d}(F_{n}-x)=2\,.$
@@
To prove that ${\rm gp}_{\rm d}(M_{k})\leq k+2$, suppose to the contrary that there exists a dual general position set $X$ of $M_{k}$ with $|X|>k+2$. We first claim that $x\notin X$ by using the argument that if $x$ would belong to $X$, then GXG-X would not be convex. So suppose that $x\in X$. Then at least one of $w_{1}$ and $w_{4}$ must belong to $X$, say $w_{1}\in X$. Then $w_{4}\notin X$. Since $w_{1}\in X$, exactly one of $w_{2}$ and $w_{k+4}$ lies in $X$. If $w_{k+4}\in X$, then $w_{2}$ and $w_{4}$ are not $X$-positionable. And if $w_{2}\in X$, then $w_{k+4}$ and $w_{4}$ are not $X$-positionable. We analogously get a contradiction in the case when $w_{4}\in X$. We can conclude that $x\notin X$. We next claim that $w_{i}\notin X$ for each $i\in[k]$. Suppose that $w_{i}\in X$ for some $i\in[k]$. Since $x\notin X$, we get that $v_{i}\in X$ because $x,w_{i},v_{i}$ is the unique shortest $x,v_{i}$-path. Further, exactly one of the neighbors of $w_{i}$ on the $(k+4)$-cycle of $W_{k+4}$ lies in $X$. But this is not possible, as $X$ is a general position set.
@@
Proof. Let $X$ be a ${\rm gp}_{\rm d}$-set of $G$ such that $x\in X$. Then $|X|={\rm gp}_{\rm d}(G)$ and by Theorem 2.2(ii), GXG-X is convex. By the proof of [7, Proposition 3.3], the set $X\setminus\{x\}$ is a general position set of GxG-x. It suffices to show that $(G-x)-(X\setminus\{x\})$ is convex in GxG-x, because then by Theorem 2.2(iii) the set $X\setminus\{x\}$ is a dual general position set of GxG-x and hence ${\rm gp}_{\rm d}(G-x)\geq|X\setminus\{x\}|={\rm gp}_{\rm d}(G)-1$.
Suppose to the contrary that this is not the case. There are $u,v\notin X$ such that a shortest $u,v$-path $P$ contains a vertex $w\in X\setminus\{x\}$. Since GXG-X is convex, $d_{G}(u,v)<d_{G}(u,w)+d_{G}(w,v)$. Based on this, we can conclude the following:
@@
This in turn implies that $x$ lies on the shortest $u,v$-path in $G$ which contradicts with GXG-X being convex. We can conclude that $(G-x)-(X\setminus\{x\})$ is convex in GxG-x.
@@
To prove that ${\rm gp}_{\rm d}(Y_{k})\leq 5$, consider an arbitrary ${\rm gp}_{\rm d}$-set $X$ of $Y_{k}$. Suppose to the contrary that $v_{i}\in X$ for some $i\in[k]$. Then exactly one of the vertices $x$ and $y$ must belong to $X$, for otherwise $v_{i}$ would lie on a shortest $x,y$-path, which is a contradiction. Without loss of generality assume that $x\in X$ and $y\notin X$. We claim that there is no vertex from $\{u,v,w\}$ which is contained in $X$. If $z\in\{u,v,w\}\cap X$, then $v_{i}$ lies on a shortest $x,z$-path, but this contradicts the fact that $X$ is a ${\rm gp}_{\rm d}$-set of $Y_{k}$. If a vertex of $\{u,v,w\}$ does not contain in $X$, say $u$, then $v_{i}$ lies on a shortest $y,u$-path, which implies that $y$ and $u$ are not $X$-positionable. This contradiction yields $v_{i}\notin X$ for each $i\in[k]$. Therefore, $|X|\leq 5$, and we conclude that ${\rm gp}_{\rm d}(Y_{k})=5$.